OpenStudy (anonymous):

A field test for a new exam was given to randomly selected seniors. The exams were graded, and the sample mean and sample standard deviation were calculated. Based on the results, the exam creator claims that on the same exam, nine times out of ten, seniors will have an average score within 5% of 75%. Is the confidence interval at 90%, 95%, or 99%? What is the margin of error? Calculate the confidence interval and explain what it means in terms of the situation. @ganeshie8

OpenStudy (anonymous):

I really need help with this problem @ganeshie8

OpenStudy (anonymous):

@ganeshie8 Please ganishie

OpenStudy (anonymous):

What does "nine times out of ten" tell you?

OpenStudy (anonymous):

The margin of error should be fairly obvious.

OpenStudy (anonymous):

As for the confidence interval, it would have the form \[\left(\bar{x}-Z_{\alpha/2}\sqrt{\frac{\sigma^2}{n}},~\bar{x}+Z_{\alpha/2}\sqrt{\frac{\sigma^2}{n}}\right)\] \(\bar{x}\) is the estimated population mean, \(Z_{\alpha/2}\) is a critical value for a \((1-\alpha)\%\) confidence level, and \(\sqrt{\dfrac{\sigma^2}{n}}\) is the margin of error. \(\bar{x}\) is given explicitly. The confidence level is just a matter of interpreting that phrase I pointed out. The margin of error is given.

OpenStudy (anonymous):

Yes it is but I somewhat need help setting this up.

OpenStudy (anonymous):

If you can help me set this up @SithsAndGiggles

OpenStudy (anonymous):

You have the formula and all the necessary info. There's not much I can do in the way of helping without just telling the answer.

OpenStudy (anonymous):

If some event occurs "nine out of ten times," what does that mean? How many times out of a hundred does the event occur?

OpenStudy (anonymous):

Ok so let me check if am actually doing this correctly 1.96√p(1-p)/N

OpenStudy (anonymous):

So if I were to plug in values it would be 1.96√9/10(1-(9/10)/?

OpenStudy (anonymous):

right @SithsAndGiggles

OpenStudy (anonymous):

Not quite, the experiment in the question is trying to estimate the population mean, not a population proportion. The 'nine times out of ten' clause should tell you something about the confidence level.

OpenStudy (anonymous):

That it would 9/10

OpenStudy (anonymous):

Right, and \(\dfrac{9}{10}=\dfrac{90}{100}=90\%\), so the interval has a 90% confidence level.

OpenStudy (anonymous):

Ok then so would I plug that into the equation

OpenStudy (anonymous):

What are the critical values for a 90% confidence interval?

OpenStudy (anonymous):

1.645

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

Correct, so \(Z_{\alpha/2}=Z_{.05}=1.645\). (Don't worry about the notation if you haven't seen it before.) The margin of error is given to be 5%, or 0.05, since an average score is expected to fall within 5% of the estimated mean. The estimated mean itself is?

OpenStudy (anonymous):

5%

OpenStudy (anonymous):

No, that's the margin of error. What's the estimated average score?

OpenStudy (anonymous):

Is it 5%+1.645? @SithsAndGiggles

OpenStudy (anonymous):

No. \[\left(\bar{x}-Z_{\alpha/2}\sqrt{\frac{\sigma^2}{n}},~\bar{x}+Z_{\alpha/2}\sqrt{\frac{\sigma^2}{n}}\right)\] \(\bar{x}\) is the estimated average score. According to the given info, that would be 75%, or 0.75. The exam creator says the scores will fall within 5% of a 75% score, so the margin of error is 5%, or 0.05. In the formula, \(\sqrt{\dfrac{\sigma^2}{n}}\) is the margin of error. The critical value for a 90% confidence interval is 1.645, so \(Z_{\alpha/2}=1.645\). So to reiterate, the confidence interval would be \[\left(0.75-1.645\times0.05,~0.75+1.645\times0.05\right)\] Simplify.

OpenStudy (anonymous):

oh ok and that would be my answer to this problem? @SithsAndGiggles

OpenStudy (anonymous):

Your answer would be the interval, yes. I suggest you study up on the topic. It doesn't seem like you're really grasping the procedure here.

OpenStudy (montverde):

Confidence Interval: 9 out of 10 = 90% The exam creator says the scores will fall within 5% of a 75% score, so the margin of error is 5%, or 0.05. The confidence interval is (.67, .83) 90% confident that the average scores fall in between 67% and 83%.

OpenStudy (montverde):

you're welcome